A stationary box of mass 42 kg is given a push of 82 N [S] along a surface where the frictional force acting is 58 N [N]. The push lasts for 0.06 minutes and then the box is allowed to slide on its own until it comes to rest. a)Draw free-body diagrams to show the box being pushed and sliding on its own., b) Determine the acceleration of the box as it is being pushed., c) Calculate the speed of the box just as the push ceases, and d)Determine the acceleration of the box as it is sliding on its own.



1
Expert's answer
2021-03-09T15:27:02-0500

a)


b)


"a=\frac{82-58}{42}=0.57\frac{m}{s^2}\ [S]"

c)


"v=0.57(0.06\cdot60)=2.05\frac{m}{s}"

d)


"a=\frac{-58}{42}=-1.38\frac{m}{s^2}\ [S]"